Circular No. 8254

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003kx AND 2003ky Further to IAUC 8252, M. Armstrong reports his discovery of two ap-

parent supernovae on unfiltered CCD images taken with a 0.35-m reflector.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003kx Dec. 8.068 8^{h}04^{m}12.^{s}35 +84^{o}38′27′′.3 16.8 18′′ W, 1′′.7 S
2003ky Dec. 8.180 11 58 06.30 +47 19 59.5 17.4 4′′.8 W, 6′′.2 S

SN 2003kx in UGC 4078 was also imaged on Dec. 9.075 UT, but nothing was visible at its location on Armstrong’s images from Oct. 17 and Nov. 6 (limiting mag 19). SN 2003ky in NGC 4001 was also imaged by Armstrong on Dec. 9.219, but nothing was visible at its location on an image from May 26 (limiting mag 19.5). Neither new object is visible on Palomar Sky Survey red and blue plates.

MARS T. Encrenaz and B. Bezard, Paris Observatory; T. K. Greathouse and

J. H. Lacy, University of Texas at Austin; M. J. Richter, University of California, Davis; and S. K. Atreya and A. S. Wong, University of Michigan, report: “On June 20, we made an unambiguous detection of hydrogen peroxide (H_{2}O_{2}) on Mars (L_{s} = 206^{o}), using high-resolution infrared spectroscopy at the NASA Infrared Telescope Facility (+ TEXES grating spectrograph) at Mauna Kea. Six individual spectral lines were identified in the range 1237–1244 cm^{−1}. The spatial distribution of H_{2}O_{2} over the Martian disk shows some enrichment in the equatorial region. The inferred H_{2}O_{2} abundance is significantly larger than the upper limit we derived from observations on 2001 Feb. 2–3 (L_{s} = 112^{o}) using the same technique (Encrenaz et al. 2002, Astron. Astrophys. 396, 1037), and it appears to be within the range of predictions from photochemical models. These observations show that H_{2}O_{2} is seasonally variable. H_{2}O_{2} has been suggested as the possible oxidant for the surface of Mars.”

CORRIGENDA On IAUC 8253, ‘Novae in M31’, line 3, the offset should read 185′′ east

and 282′′ south

On IAUC 8236, ‘Supernova 2003jg in NGC 2997’, line 3, for SN 2000jg read SN 2003jg

2003 December 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green